OpenStudy (bananas1234):

Which expression makes the equation true? 64h^8 = (blank)^2 32h^8 32h^4 8h^4 8h^8

OpenStudy (bananas1234):

would this be 32h^8? @rational

OpenStudy (rational):

put that in the blank and see if you get back the left hand side expression\[\left(\text{blank}\right)^2\]

OpenStudy (rational):

\[\left(32h^8\right)^2 = ?\]

OpenStudy (bananas1234):

hmmm, sorry i am confused on how to solve (32^8)^2 :/

OpenStudy (bananas1234):

1.2089258e+24?